In 2015, the mean return of all common stocks on the Philippine Stock Exchange was 0.45. The standard deviation of the returns was about 6.2. A student of finance forms all possible portfolios that invested equal amounts in 4 of these stocks and records the return for each portfolio. This return is the average of the returns of the 4 stocks chosen. What is the variance of the portfolio returns?


1
Expert's answer
2022-03-07T08:04:03-0500

Let X be a random variable representing the average return of the 4 stocks, Y - representing return of 1 stock, then "V(Y)={\frac {V(X)} 4}={\frac {6.2^2} 4}={\frac {38.44} 4}=9.61"